In the lab, Ivan has two solutions that contain alcohol and is mixing them with each other. He uses 300 milliliters less of Solution A than Solution B. Solution A is 11% alcohol and Solution B is 16% alcohol. How many milliliters of Solution B does he use, if the resulting mixture has 291 milliliters of pure alcohol?


Answer by stanbon(75887)

He uses 300 milliliters less of Solution A than Solution B. Solution A is 11% alcohol and Solution B is 16% alcohol. How many milliliters of Solution B does he use, if the resulting mixture has 291 milliliters of pure alcohol?
-----
Equation:
alcohol + alcohol = alcohol
A = B-300 ml
---
0.11A + 0.16B = 291
---
0.11(B-300) + 0.16B = 291
Multiply thru by 100 to get:
11B - 3300 + 16B = 29100
27B = 32400
B = 1200 ml (amt. of solution B needed)
A = B-300 = 900 ml (amt. of solution A needed
===============
Cheers,
Stan H.
===============